# Constructing new probability distributions on the unit interval

### v1

@@ 8 Conclusions
We have introduced two flexible families of unit‐interval distributions by applying non‐injective transformations to the ratio of independent gamma variates, obtaining closed‐form expressions for their densities and distribution functions. We have derived quantile representations, moments in special cases, and established a maximum likelihood estimation procedure for joint inference on $(\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2},r)$. As part of future research, it will be of interest to explore regression extensions, Bayesian methods, and applications to bounded‐response data.
@@ Acknowledgements
Appendix A Appendix
@@
$\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]$ $u\exp(-ux){\rm d}F_{X}(u)$
$={\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}},$
@@
[…]
@@
$\int_{0}^{\infty}{1\over(ax+b)^{p}(cx+e)^{q}}{\rm d}x$ $={a^{q}\over ab^{p-1}c^{q}b^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}({bc\over a}y+e)^{q}}{\rm d}y$
$={a^{q-1}\over b^{p+q-1}c^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}(y+{ae\over bc})^{q}}{\rm d}y.$ (37)
@@
with $a=q$, $b=p+q-1$, $c=p+q$ and $z=1-ae/cb$, expression in (A) becomes
@@
$\xi(p)={1\over p+1}\Bigg{\{}{\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]\\[5.69046pt] +{\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1)-\left({1\over 2}\right)^{p+1}\Bigg{\}},$
@@
$={1\over p+1}\left\{\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]-\left({1\over 2}\right)^{p+1}\mathbb{P}\left({X\over X+Y}>{1\over 2}\right)\right.$
@@
$\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]$ $=\int_{1\over 2}^{1}\frac{z^{(\alpha_{1}+p+1)-1}(1-z)^{\alpha_{2}-1}}{\mathrm{B}(\alpha_{1},\alpha_{2})}{\rm d}z$
$={\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\int_{1\over 2}^{1}\frac{z^{(\alpha_{1}+p+1)-1}(1-z)^{\alpha_{2}-1}}{\mathrm{B}(\alpha_{1}+k+1,\alpha_{2})}{\rm d}z$
$={\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]$ (40)
@@
$\mathbb{E}\left[\left({Y\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{Y\over X+Y}<1\}}\right]$ $={\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{2}+p+1,\alpha_{1})\right]$
$={\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1),$ (41)
@@
By plugging (A), (A) and (42) in (39), the proof follows. ∎

### v2

@@ 10 Conclusions
This work introduces two novel families of probability distributions on the unit interval, obtained through non-injective transformations of the gamma ratio. The proposed models extend classical unit-interval distributions and establish a direct connection with well-known inequality measures, such as the Gini and Atkinson indices. The derived closed-form expressions for density, cumulative distribution, and raw moments, alongside maximum likelihood estimation procedures, demonstrate both analytical tractability and practical applicability. The flexibility of the models is highlighted through the parameter $r$, which controls the shape of the resulting distributions, allowing the capture of diverse features such as unimodality and skewness. Potential applications span econometrics, reliability analysis, and other fields where modeling proportions or normalized quantities is essential. Future work may explore correlated gamma components, extend the transformations to multivariate settings, and investigate Bayesian estimation procedures for these models.
@@ Acknowledgements
Appendix A Some technical results
@@
$\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)={\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}},$
@@
where $_{2}F_{1}(a,b;c;z)$ is the hypergeometric function, defined for $|z|<1$ by the power series
@@
$\int_{0}^{\infty}{1\over(ax+b)^{p}(cx+e)^{q}}{\rm d}x$ $={a^{q}\over ab^{p-1}c^{q}b^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}({bc\over a}y+e)^{q}}{\rm d}y{a^{q-1}\over b^{p+q-1}c^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}(y+{ae\over bc})^{q}}{\rm d}y.$ (35)
@@
with $a=q$, b=p+q1b=p+q-1, $c=p+q$ and z=1ae/cbz=1-ae/cb, expression in (35) becomes
@@
$\xi(p)={1\over p+1}\Bigg\{{\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]+{\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1)-\left({1\over 2}\right)^{p+1}\Bigg\},$
@@
${1\over p+1}\left\{\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]-\left({1\over 2}\right)^{p+1}\mathbb{P}\left({X\over X+Y}>{1\over 2}\right)\right$.
@@
$\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]=\int_{1\over 2}^{1}\frac{z^{(\alpha_{1}+p+1)-1}(1-z)^{\alpha_{2}-1}}{\mathrm{B}(\alpha_{1},\alpha_{2})}{\rm d}z={\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]$ (38)
@@
$\mathbb{E}\left[\left({Y\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{Y\over X+Y}<1\}}\right]={\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1),$ (39)
@@
By plugging (38), (39) and (40) in (37), the proof follows. ∎
Appendix B Complete results of the simulation study
This appendix gathers additional results from the simulation study, with particular emphasis on the scenarios in which the scale parameters are kept fixed ($\lambda_{1}=\lambda_{2}=1$). This configuration makes it possible to assess separately the impact of the shape parameters $(\alpha_{1},\alpha_{2})$ and of the transformation parameter $r$ on estimator performance, isolating the effect of scale. For these models, the bias and RMSE panels obtained from 500 Monte Carlo replications are reported for each shape scenario, considering $n\in\{50,100,250,500\}$ and $r\in\{0{.}5,1,2,4\}$. The results corresponding to the full estimation of the five parameters $(\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2},r)$ are not reproduced here, since the main interpretation—based exclusively on the bias and RMSE plots—is fully presented in the main text.
B.1 Model $W$ — Estimation with $\lambda_{1}=\lambda_{2}=1$
This section presents the detailed simulation results for model $W$ under $\lambda_{1}=\lambda_{2}=1$. In this setting, only the shape parameters $(\alpha_{1},\alpha_{2})$ and the transformation parameter $r$ are estimated by maximum likelihood. The analysis focuses on the behavior of the estimators across different sample sizes and values of $r$, highlighting patterns of bias and variability.
The scenarios considered are $(\alpha_{1},\alpha_{2})\in\{(0{.}5,0{.}5),\ (0{.}8,1{.}0),\ (1{.}5,1{.}3),\ (2{.}0,2{.}0)\}.$
B.1.1 Scenario $(\alpha_{1},\alpha_{2})=(0{.}5,0{.}5)$
Figure 17 presents the bias and RMSE plots for the (0.5,0.5)(0{.}5,0{.}5) scenario. This is a low-shape regime, characterized by a strong concentration of mass near zero. As a consequence, the estimates exhibit higher variability, especially for the transformation parameter $r$.
The estimates of $\alpha_{1}$ and $\alpha_{2}$ display small biases and RMSE values around 0.090{.}090.120{.}12 at $n=50$, with a consistent reduction as $n$ increases. In contrast, the errors associated with $\widehat{r}$ increase rapidly with $r$, reaching high levels for $r=2$ and $r=4$ in small samples. Although increasing the sample size substantially reduces these errors, the estimation of $r$ remains more unstable than that of the shape parameters.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 17: Bias and RMSE of the estimators for model $W$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(0.5,0.5)$.
B.1.2 Scenario $(\alpha_{1},\alpha_{2})=(0{.}8,1{.}0)$
The (0.8,1.0)(0{.}8,1{.}0) scenario exhibits moderate asymmetry, leading to improved stability relative to the (0.5,0.5)(0{.}5,0{.}5) case. The plots in Figure 18 show a gradual reduction in bias and RMSE for the shape parameters, especially from $n=100$ onward.
Even so, the parameter $r$ remains highly sensitive in small samples: errors increase with larger values of $r$, but decrease substantially at $n=250$ and stabilize at $n=500$. This behavior reinforces that the estimation of $r$ tends to require larger sample sizes to achieve precision comparable to that obtained for $(\alpha_{1},\alpha_{2})$.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 18: Bias and RMSE of the estimators for model $W$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(0.8,1.0)$.
B.1.3 Scenario $(\alpha_{1},\alpha_{2})=(1{.}5,1{.}3)$
The (1.5,1.3)(1{.}5,1{.}3) scenario combines moderate asymmetry with higher shape values, yielding noticeably more stable estimators. Figure 19 shows that the shape parameters have small biases and RMSE values that decrease consistently as $n$ grows.
Despite the overall improvement, $\widehat{r}$ still concentrates the largest variability in the model, particularly for $r=4$. The pattern, however, is smoother than in low-shape scenarios, indicating that larger values of $(\alpha_{1},\alpha_{2})$ help stabilize the estimation procedure.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 19: Bias and RMSE of the estimators for model $W$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(1.5,1.3)$.
B.1.4 Scenario $(\alpha_{1},\alpha_{2})=(2{.}0,2{.}0)$
In the symmetric (2.0,2.0)(2{.}0,2{.}0) scenario, the most stable behavior among the four cases is observed. Figure 20 shows that the shape parameters exhibit small biases and RMSE values that decrease rapidly as $n$ increases. Symmetry contributes to the regularity of the estimates and to the homogeneity of errors between $\widehat{\alpha}_{1}$ and $\widehat{\alpha}_{2}$.
Although $\widehat{r}$ still shows sensitivity to larger values of $r$, its variability decreases more markedly than in the previous scenarios, reinforcing the stabilizing effect of symmetry and higher shape values.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 20: Bias and RMSE of the estimators for model $W$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(2.0,2.0)$.
B.2 Model $Z$ — Estimation with $\lambda_{1}=\lambda_{2}=1$
This appendix presents the complete results of the simulation study for model $Z$ under the configuration in which the scale parameters are kept fixed, with $\lambda_{1}=\lambda_{2}=1$. Unlike the full estimation approach, in this configuration only the shape parameters $(\alpha_{1},\alpha_{2})$ and the transformation parameter $r$ are estimated.
The results show that, in contrast to model $W$, fixing the scale parameters is not sufficient to stabilize the estimators for model $Z$. This instability manifests through large biases and high RMSE values, revealing important challenges in parameter identifiability even at large sample sizes.
B.2.1 Scenario $(\alpha_{1},\alpha_{2})=(0{.}5,0{.}5)$
In the symmetric low-shape scenario $(\alpha_{1},\alpha_{2})=(0{.}5,0{.}5)$, the estimators $\widehat{\alpha}_{1}$ and $\widehat{\alpha}_{2}$ exhibit strong overestimation for all combinations of $n$ and $r$. At $n=50$, biases exceed $15$ for $r\leq 2$, with RMSE values above $50$. Although there is a progressive reduction as $n$ increases, even at $n=500$ the RMSE remains between $7$ and $30$ for $r\leq 2$, indicating substantial identifiability difficulties. Only for $r=4$ is a marked improvement observed, with nearly zero bias and RMSE around 0.20{.}2 at $n=500$.
The transformation parameter $\widehat{r}$ is even more unstable. For $r\in\{0{.}5,1,2\}$, the biases at $n=50$ range from 21.921{.}9 to 81.481{.}4, decreasing slowly to values between 2.82{.}8 and 11.711{.}7 at $n=500$. The RMSE remains very large across the entire sample-size range (above $40$ for $r\leq 2$). Only for $r=4$ is satisfactory convergence observed, with RMSE close to 1.81{.}8 at $n=500$.
This scenario is one of the most challenging regimes for model $Z$, due to the strong concentration of probability mass near the boundaries when $\alpha_{1}=\alpha_{2}<1$.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 21: Bias and RMSE of the estimators for model $Z$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(0.5,0.5)$.
B.2.2 Scenario $(\alpha_{1},\alpha_{2})=(0{.}8,1{.}0)$
In the moderately asymmetric scenario, the shape estimators remain overestimated for all combinations of $n$ and $r$. At $n=50$, biases range between $12$ and $21$, with RMSE values exceeding $45$. Although errors decrease gradually as $n$ increases, biases between $3$ and $7$ and RMSE values above $10$ are still observed at $n=500$. Asymmetry has a clear impact: larger values of $r$ amplify biases and RMSE values, especially for $\widehat{\alpha}_{1}$.
The estimation of $r$ in this scenario is highly unstable for all true parameter values. At $n=50$, biases range from about 77 (r=0.5r=0{.}5) to more than $65$ ($r=4$), with RMSE between $21$ and $177$. Increasing $n$ partially reduces errors, but they remain high even at $n=500$, especially for $r=2$ and $r=4$, whose RMSE values remain close to $60$. Thus, the identification of $r$ remains quite problematic in this scenario.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 22: Bias and RMSE of the estimators for model $Z$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(0.8,1.0)$.
B.2.3 Scenario $(\alpha_{1},\alpha_{2})=(1{.}5,1{.}3)$
In contrast to the previous two scenarios, the (1.5,1.3)(1{.}5,1{.}3) case exhibits much more favorable performance. The shape estimators show moderate biases that decrease as $n$ increases. At $n=50$, the bias of $\widehat{\alpha}_{1}$ ranges between 0.130{.}13 and 0.160{.}16, whereas the bias of $\widehat{\alpha}_{2}$ lies between 0.200{.}20 and 0.320{.}32. The RMSE values decrease from approximately 0.40{.}41.01{.}0 to ranges between 0.20{.}2 and 0.40{.}4 when $n$ reaches $500$, indicating adequate precision and little impact of the value of $r$.
The transformation parameter is recovered very well for $r\leq 2$, with biases close to zero and RMSE below 0.100{.}10 for $n\geq 250$. For $r=4$, although moderate negative biases persist, the RMSE values remain relatively small (between 0.190{.}19 and 0.210{.}21 at $n=500$). This scenario shows that model $Z$ can be estimated with good stability when the generating distribution has moderate shape and is not excessively asymmetric.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 23: Bias and RMSE of the estimators for model $Z$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(1.5,1.3)$.
B.2.4 Scenario $(\alpha_{1},\alpha_{2})=(2{.}0,2{.}0)$
Despite symmetry and higher shape, the (2.0,2.0)(2{.}0,2{.}0) scenario shows unsatisfactory performance for model $Z$ with fixed scales. The shape estimators remain strongly overestimated across all sample sizes, with biases ranging between $20$ and $40$ and RMSE values often above $70$, even at $n=500$. Increases in $n$ lead only to modest reductions.
The estimation of $r$ is also problematic: for $r\leq 2$, RMSE values range from $13$ to $98$, and for $r=4$ they reach values close to $200$, indicating severe instability. These results suggest that the combination of larger shape values and the transformation $T_{r}(x)$ produces densities that contain relatively little information for estimating the parameters of model $Z$, even with $\lambda_{1}$ and $\lambda_{2}$ fixed.
(a) Bias of the estimators
(b) RMSE of the estimators
Figure 24: Bias and RMSE of the estimators for model $Z$ with $\lambda_{1}=\lambda_{2}=1$ under the scenario $(\alpha_{1},\alpha_{2})=(2.0,2.0)$.
Unlike what is observed for model $W$, fixing the scale parameters is not sufficient to stabilize estimation in model $Z$, especially in low-shape or asymmetric scenarios. In these cases, errors remain large even at $n=500$, particularly for the transformation parameter. Only in the moderate-shape scenario (1.5,1.3)(1{.}5,1{.}3) is satisfactory performance observed. These results reinforce the need for larger sample sizes or alternative estimation methods for model $Z$.